# Teaching MLPs to Master Heterogeneous Graph-Structured Knowledge for Efficient and Accurate Inference

### v1

@@ I Introduction
Recently, heterogeneous graph neural networks (HGNNs) have gained recognition for their ability to capture the intricate heterogeneity and rich semantic information inherent in heterogeneous graph structures [7, 8, 9, 10, 11, 12]. HGNNs generate node representations through recursive aggregation of messages from diverse types of neighboring nodes across various edge types. By adhering to this heterogeneity-aware message-passing paradigm, HGNNs adeptly model the complex relationships, diverse entity types, and heterogeneous semantics present in heterogeneous graphs, achieving state-of-the-art performance in various downstream tasks, such as node classification [13], node clustering [14], community search [15], and link prediction [16].
@@
While HGNNs have achieved notable advancements, deploying them for real-world industrial applications remains challenging, especially in environments characterized by large-scale data, limited memory, and high sensitivity to latency. The primary obstacle lies in HGNNs’ dependence on the heterogeneous graph structure during inference. Specifically, inference on a target node necessitates fetching the features of numerous neighboring nodes according to the graph topology, causing the number of nodes fetched and inference times to grow exponentially with the number of HGNN layers [18]. Additionally, HGNNs process different edge types independently and require distinct mapping functions for each node type, resulting in increased computational overhead. Figure 1 shows that increasing the number of HGNN layers intensifies node dependencies, exponentially boosting the count of fetched nodes and inference time. Conversely, MLPs exhibit much smaller and linearly growing inference time as they only take the target node feature as input. However, this lack of graph dependency typically limits MLPs’ performance on downstream tasks compared to HGNNs.
@@ III-A Heterogeneous Graph
As shown in Figure 2(b), two movies can be connected via two meta-paths: Movie-Actor-Movie (MAM) and Movie-Director-Movie (MDM). Meta-path usually reveals heterogeneous semantics in a heterogeneous graph. For example, MAM means the co-actor relation, while MDM indicates that two movies are directed by the same director.
@@ IV-B HG2M+
Despite the effectiveness of HG2M, it faces two critical challenges: (1) The soft labels of nodes incorrectly predicted by teacher HGNNs will introduce noises to the optimization landscape of student MLPs. (2) HG2M ignores the intricate relationships and diverse relational semantics inherent in heterogeneous graph structures. Therefore, we further propose HG2M+ to distill reliable and heterogeneous semantic knowledge into student MLPs through reliable node distillation and reliable meta-path distillation, as shown in Figure 3.
@@ IV-B2 Reliable Meta-Path Distillation
$\mathcal{L}=\frac{\lambda}{|\mathcal{V}_{t}^{L}|}\sum_{v\in\mathcal{V}_{t}^{L}}\mathcal{L}_{CE}\left(\hat{\boldsymbol{y}}_{v},\boldsymbol{y}_{v}\right)+(1-\lambda)\left(\mathcal{L}_{RND}+\mathcal{L}_{RMPD}\right).$ (6)
@@ IV-C Why do HG2Ms work?
where $I(\mathcal{E}^{[v]};\boldsymbol{y}_{v})$ is the mutual information between edges and labels, indicating the relevance of the graph structure to labels, while $I(\boldsymbol{X}^{[v]};\boldsymbol{y}_{v}|\mathcal{E}^{[v]})$ is the mutual information between features and labels conditioned on edges $\mathcal{E}^{[v]}$. HG2M leverages the objective function defined in Eq. (2), which approximates $I(\mathcal{G}^{[v]};\boldsymbol{y}_{v})$ by only maxmizing $I(\boldsymbol{X}^{[v]};\boldsymbol{y}_{v}|\mathcal{E}^{[v]})$ while ignoring $I(\mathcal{E}^{[v]};\boldsymbol{y}_{v})$. But for real-world node classification tasks, node features and structural roles are often highly correlated [37, 18], allowing MLPs to perform reasonably well even when relying solely on node features. Then with the help of knowledge distillation, MLP parameters approximate the ideal prediction function from node features to labels, thus HG2M can potentially achieve better results. Additionally, to leverage the complementary structural information, HG2M+ uses meta-path structures to model $I(\mathcal{E}^{[v]};\boldsymbol{y}_{v})$, so it can work better in the case that labels are also correlated to the graph structure. Even in the extreme case where $\boldsymbol{y}_{v}$ is uncorrelated with $I(\boldsymbol{X}^{[v]};\boldsymbol{y}_{v}|\mathcal{E}^{[v]})$, HG2M+ can still achieve superior or comparable performance to HGNNs, as shown in Section V-G.
@@ V Experiments
How does each component contribute to in HG2M+?
@@ V-A1 Datasets
We evaluate HG2Ms on six real-world datasets, including TMDB, CroVal, ArXiv, and three more larger IGB datasets [17]: IGB-549K-19, IGB-549K-2K, and IGB-3M-19. In Table I, we present the statistical information of the datasets used in our experiments. Further details about each dataset are provided below.
@@ V-A2 Model Architectures
For a fair comparison, we use RSAGE, a GraphSAGE [41] extension to heterogeneous graphs, with GCN [23] aggregation as the teacher. We also investigate the influence of alternative teacher models, including RGCN [7], RGAT [42], SimpleHGN [11], and ieHGCN [10], as detailed in Section V-F.
@@ V-A4 Implementation Details
The experiments on both baselines and our approach are implemented using PyTorch and DGL for GNN algorithms, and Adam [43] for optimization. We run all experiments on a 32GB NVIDIA Tesla V100 GPU. The hyperparameters for vanilla MLPs, teacher HGNNs, and HG2Ms are detailed in Table II. HG2Ms are configured with the same number of layers and hidden dimensions per layer as teacher HGNNs to maintain the same parameter counts. The parameter $\lambda$ in Eq. (2) and Eq. (6) balances the weight of supervision from the true labels and soft labels. In our experiments, we performed slight tuning of $\lambda$ but found that non-zero values did not significantly improve results. Therefore, following the approach in [18, 29], we present results with $\lambda=0$, where only the soft label term is effective. Additionally, based on empirical findings in Section V-H2, we set the reliable node proportion $p=0.9$ consistently across all datasets. For RMPD in HG2M+, we use the meta-paths listed in Table I. To ensure reproducibility, our implementation is publically available at https://github.com/Cloudy1225/HG2M.
@@ V-C How do HG2Ms compare to other GNNs-to-MLPs?(RQ2)
Heterogeneous graphs present more complex semantic relationships between nodes, requiring models that effectively capture such heterogeneous semantics. While previous GNN-to-MLP methods have mainly focused on homogeneous graphs, applying them to heterogeneous graphs may lead to suboptimal performance. Thus, we meticulously design HG2M+, which integrates RND and RMPD, to distill reliable and heterogeneous knowledge from HGNNs to MLPs.
TABLE V: Transductive accuracy of different GNN-to-MLP methods.
@@
To validate the effectiveness of HG2M+, we further compare it with four previous GNN-to-MLP knowledge distillation frameworks: RKD-MLP [25], FF-G2M [26], KRD [20], and NOSMOG [19], which leverage more complex techniques like feature-based distillation and representational similarity distillation. Table V presents the transductive accuracy of these models on the TMDB, CroVal, and IGB-549K-19 datasets. We can find that HG2M+ consistently achieves the best performance. This highlights the superiority of our HG2M+ in the context of heterogeneous graphs.
@@ V-E How does each component contribute to HG2M+? (RQ4)
Since HG2M+ contains two essential components (i.e., reliable node distillation (RND) and reliable meta-path distillation (RMPD)), we conduct ablation studies to assess the individual contributions of each component when integrated independently into HG2M. Table VII demonstrates that performance improves with the addition of each component, indicating their effectiveness. In general, the incorporation of RND contributes a lot, as RND can filter out noisy information, making heterogeneous semantic knowledge distilled by RMPD more reliable.
TABLE VII: Ablation study for independent components in HG2M+. The increasing performance shows that each component in HG2M+ contributes positively to the KD process.
@@
Figure 4: Effectiveness of our reliable node selection strategy in RND and intra-class meta-path-based neighbor estimation strategy in RMPD. (a) The accuracy of selected reliable nodes $\mathcal{R}^{U}$ (orange) is much higher than that of the raw unlabeled node set $\mathcal{V}_{t}^{U}$ (blue). (b) The homophily of selected meta-path-based neighbor pairs $\mathcal{M}_{P}^{U}$ (orange) is much higher than that of the raw meta-path-based neighbor pair set $\mathcal{E}_{P}$ (blue).
@@
We also conduct an in-depth analysis to demonstrate the effectiveness of our reliable node selection strategy in RND and our intra-class meta-path-based neighbor estimation strategy in RMPD. Figure 4(a) presents the accuracy of raw unlabelled nodes $\mathcal{V}_{t}^{U}$, selected unlabelled reliable nodes $\mathcal{R}^{U}$, and all selected reliable nodes $\mathcal{R}$ with a selection proportion of $p=0.9$. Figure 4(b) displays the homophily of all meta-path-$P$-based neighbor pairs $\mathcal{E}_{P}$, selected unlabelled meta-path-$P$-based neighbor pairs $\mathcal{M}_{P}^{U}$, and all selected meta-path-$P$-based neighbor pairs $\mathcal{M}_{P}$. Here, homophily represents the fraction of intra-class neighbor pairs, with meta-paths MAM, QTQ, PAP, and PAP applied to TMDB, CroVal, ArXiv, and IGBs respectively. We observed that the accuracy of selected unlabelled reliable nodes (orange) is significantly higher than that of raw unlabelled reliable nodes (blue), and the homophily of selected unlabelled meta-path-$P$-based neighbor pairs (orange) is also markedly higher than that of all meta-path-$P$-based neighbor pairs (blue). Furthermore, the accuracy of the final reliable nodes (green) used for distillation and the homophily of the final reliable intra-class meta-path-based neighbor pairs (green) used for distillation both exceed 0.9. These findings underscore the efficacy of our reliable node selection strategy in RND for identifying reliable nodes and our intra-class meta-path-based neighbor estimation strategy in RMPD for identifying intra-class neighbors.
V-F How do HG2Ms perform with different teachers? (RQ5)
@@
V-G How do HG2Ms perform with noisy node features? (RQ6)
Considering that HG2Ms utilize only node features as input, they may be susceptible to feature noise and could underperform when the labels are unrelated to the node features. Therefore, we evaluate the robustness of HG2Ms with regards to different noise levels across TMDB, CroVal, and IGB-549K-19 datasets. Specifically, we introduce different levels of Gaussian noises to node features by replacing $\boldsymbol{X}_{t}$ with $(1-\alpha)\boldsymbol{X}_{t}+\alpha\epsilon$, where $\epsilon$ is independent isotropic Gaussian noise, and $\alpha\in[0,1]$ controls the noise level. Our results, depicted in Figure 6, reveal that HG2Ms achieve comparable or improved performance compared to teacher HGNNs across different noise levels. Moreover, HG2Ms not only outperform MLPs but also exhibit slower performance degradation as $\alpha$ increases. When $\alpha$ approaches 1, the input features and node labels will become independent corresponding to the extreme case discussed in Section IV-C. In this scenario, HG2M+ continues to perform as well as HGNNs, while MLPs perform poorly.
V-H How do different hyperparameters affect HG2Ms? (RQ7)
V-H1 Inductive Split Rate

### v2

@@ I Introduction
Recently, heterogeneous graph neural networks (HGNNs) have gained recognition for their ability to capture heterogeneous semantic knowledge that emerges from multiple node and edge types, as well as complex relational patterns [7, 8, 9, 10]. In particular, HGNNs effectively model relational semantics, which describe the meanings of different types of connections between entities (e.g., co-actor or co-director relationships in Figure 2) [11, 12]. HGNNs learn node representations by recursively aggregating information from neighboring nodes of various types across different relations. Through this heterogeneity-aware message-passing mechanism, they can model complex interactions, diverse entity types, and rich semantic structures, achieving state-of-the-art results in tasks such as node classification [13], node clustering [14], and link prediction [15].
@@
While HGNNs have achieved notable advancement, deploying them in real-world industrial settings remains challenging, especially in scenarios involving large-scale data, limited memory, and strict latency requirements. A key limitation is their reliance on the graph structure during inference. For each target node, HGNNs must fetch features from many neighbors based on the graph topology, causing both the number of fetched nodes and the inference time to grow exponentially with the number of layers [17]. Moreover, HGNNs process each edge type separately and often require distinct mapping functions for different node types, further increasing computational overhead. As illustrated in Figure 1, deeper HGNNs result in significantly higher node dependencies and inference latency. In contrast, MLPs perform inference using only the target node’s features, leading to much lower and linearly scaling inference time. However, this independence from graph structure often limits MLPs’ effectiveness on downstream tasks compared to HGNNs.
@@ III-A Heterogeneous Graph
As shown in Figure 2(b), two movies can be connected via two meta-paths: Movie-Actor-Movie (MAM) and Movie-Director-Movie (MDM). Meta-path usually reveals heterogeneous semantics in a heterogeneous graph [11, 12, 13]. For example, MAM means the co-actor relation, while MDM indicates that two movies are directed by the same director.
@@ IV-B HG2M+
Despite the effectiveness of HG2M, it faces two critical challenges: (1) The soft labels of nodes incorrectly predicted by teacher HGNNs introduce noise into student MLP training, which can hinder generalization. (2) HG2M overlooks the intricate and diverse semantic relationships embedded in heterogeneous graph structures, as it only distills node-level logits without considering relational context. Therefore, we further propose HG2M+, which incorporates Reliable Node Distillation (RND) to filter out noisy labels by selecting only high-confidence, low-uncertainty predictions [29] from the teacher HGNNs, and Reliable Meta-path Distillation (RMPD) to capture higher-order heterogeneous semantics by leveraging reliable, intra-class meta-path-based neighbors [11, 12, 13, 15]. These enhancements allow HG2M+ to better utilize both reliable supervision and rich heterogeneous structure, improving over the HG2M in both accuracy and robustness. The framework of HG2M+ is presented in Figure 3.
@@ IV-B2 Reliable Meta-Path Distillation
$\mathcal{L}$ $=\frac{\lambda}{|\mathcal{V}_{t}^{L}|}\sum_{v\in\mathcal{V}_{t}^{L}}\mathcal{L}_{CE}\left(\hat{\boldsymbol{y}}_{v},\boldsymbol{y}_{v}\right)+(1-\lambda)(\frac{1}{|\mathcal{R}|}\sum_{v\in\mathcal{R}}\mathcal{L}_{KL}\left(\hat{\boldsymbol{y}}_{v},\boldsymbol{z}_{v}\right)$ (6)
$+\frac{1}{|\mathcal{P}|}\sum_{P\in\mathcal{P}}\frac{1}{|\mathcal{M}_{P}|}\sum_{(u,v)\in\mathcal{M}_{P}}p_{uv}\mathcal{L}_{KL}(\hat{\boldsymbol{y}}_{v},\boldsymbol{z}_{u})\Large).$
@@ IV-C Why do HG2Ms work?
where $I(\mathcal{E}^{[v]};\boldsymbol{y}_{v})$ is the mutual information between edges and labels, indicating the relevance of the graph structure to labels, while $I(\boldsymbol{X}^{[v]};\boldsymbol{y}_{v}|\mathcal{E}^{[v]})$ is the mutual information between features and labels conditioned on edges $\mathcal{E}^{[v]}$. HG2M leverages the objective function defined in Eq. (2), which approximates $I(\mathcal{G}^{[v]};\boldsymbol{y}_{v})$ by only maxmizing $I(\boldsymbol{X}^{[v]};\boldsymbol{y}_{v}|\mathcal{E}^{[v]})$ while ignoring $I(\mathcal{E}^{[v]};\boldsymbol{y}_{v})$. But for real-world node classification tasks, node features and structural roles are often highly correlated [36, 17], allowing MLPs to perform reasonably well even when relying solely on node features. Then with the help of knowledge distillation, MLP parameters approximate the ideal prediction function from node features to labels, thus HG2M can potentially achieve better results. Additionally, to leverage the complementary structural information, HG2M+ uses meta-path structures to model $I(\mathcal{E}^{[v]};\boldsymbol{y}_{v})$, so it can work better in the case that labels are also correlated to the graph structure. Even in the extreme case where $\boldsymbol{y}_{v}$ is uncorrelated with $I(\boldsymbol{X}^{[v]};\boldsymbol{y}_{v}|\mathcal{E}^{[v]})$, HG2M+ can still achieve superior or comparable performance to HGNNs, as shown in Section V-H.
@@ V Experiments
How does each component contribute to HG2M+?
@@ V-A1 Datasets
We evaluate HG2Ms on six real-world datasets, including three small datasets TMDB, CroVal, ArXiv [37] and three more larger IGB datasets [16]: IGB-549K-19, IGB-549K-2K, and IGB-3M-19. In Table I, we present the statistical information of the datasets used in our experiments. Further details about each dataset are provided below.
@@ V-A2 Model Architectures
For a fair comparison, we use RSAGE, a GraphSAGE [41] extension to heterogeneous graphs, with GCN [22] aggregation as the teacher. We also investigate the influence of alternative teacher models, including RGCN [7], RGAT [42], SimpleHGN [9], and ieHGCN [8], as detailed in Section V-G.
@@ V-A4 Implementation Details
The experiments on both baselines and our approach are implemented using PyTorch and DGL for GNN algorithms, and Adam [43] for optimization. We run all experiments on a 32GB NVIDIA Tesla V100 GPU. The hyperparameters for vanilla MLPs, teacher HGNNs, and HG2Ms are detailed in Table II. HG2Ms are configured with the same number of layers and hidden dimensions per layer as teacher HGNNs to maintain the same parameter counts. The parameter $\lambda$ in Eq. (2) and Eq. (6) balances the weight of supervision from the true labels and soft labels. In our experiments, we performed slight tuning of $\lambda$ but found that non-zero values did not significantly improve results. Therefore, following the approach in [17, 28, 44], we present results with $\lambda=0$, where only the distillation term is effective. Additionally, based on empirical findings in Section V-I2, we set the reliable node proportion $p=0.9$ consistently across all datasets. For RMPD in HG2M+, we use the meta-paths listed in Table I. To ensure reproducibility, our implementation is publicly available at https://github.com/Cloudy1225/HG2M.
@@ V-C How do HG2Ms compare to other GNNs-to-MLPs?(RQ2)
While previous GNN-to-MLP methods have mainly focused on homogeneous graphs, applying them to heterogeneous graphs may lead to suboptimal performance, as they fail to consider the intrinsic heterogeneity and semantic richness captured by meta-path structures. Thus, we meticulously design HG2M+, which integrates RND and RMPD, to distill reliable and heterogeneous knowledge from HGNNs to MLPs.
TABLE V: Transductive accuracy of different distillation methods.
@@
To validate the effectiveness of HG2M+, we further compare it with seven previous knowledge distillation frameworks: RKD-MLP [24], FF-G2M [25], KRD [19], NOSMOG [18], HIRE [45], TGS [46], and TeKAP [47], which leverage more complex techniques like ensemble distillation, feature-based distillation, and representational similarity distillation. Table V presents the performance of these models on the TMDB, CroVal, and IGB-549K-19 datasets. We can find that HG2M+ consistently achieves the best performance. This highlights the superiority of our HG2M+ for heterogeneous graphs.
@@ V-D How efficient are HG2Ms compared to HGNNs? (RQ3)
We further present an extended efficiency analysis on the largest dataset IGB-3M-19, and report the time cost and GPU memory usage (in MB) for both training and inference phases. The inference results are also based on 5 randomly selected test nodes. As shown in Table VII, despite incurring slightly higher training time than RSAGE (within 5%), HG2Ms require less than half the training memory. Most notably, at inference time, HG2Ms achieve 379× speedup and 28× lower memory usage while maintaining competitive or better accuracy than RSAGE. These results further highlight the practical advantages of our framework for deployment in large-scale and memory-constrained settings.
TABLE VII: Comprehensive efficiency analysis on IGB-3M-19.
Method Train Time (s) Train Mem. Infer. Time (ms) Infer. Mem.
MLP 455.62 5401.76 0.33 9.66
RSAGE 3500.21 23784.22 125.15 275.70
HG2M 3615.88 (0.97×) 9650.93 (2.46×) 0.33 (379.24×) 9.66 (28.54×)
HG2M+ 3698.23 (0.95×) 11591.48 (2.05×) 0.33 (379.24×) 9.66 (28.54×)
@@ V-E How does each component contribute to HG2M+? (RQ4)
Since HG2M+ contains two essential components (i.e., reliable node distillation (RND) and reliable meta-path distillation (RMPD)), we conduct ablation studies to assess the individual contributions of each component when integrated independently into HG2M. Table VIII demonstrates that performance improves with the addition of each component, indicating their effectiveness. In general, the incorporation of RND contributes a lot, as RND can filter out noisy information, making the heterogeneous semantic knowledge distilled by RMPD more reliable.
TABLE VIII: Ablation study for independent components in HG2M+. The increasing performance shows that each component in HG2M+ contributes positively to the KD process.
@@
Figure 4: Effectiveness of our reliable node selection in RND and intra-class meta-path-based neighbor estimation in RMPD. (a) The accuracy of selected reliable nodes $\mathcal{R}^{U}$ (orange) is much higher than that of the raw unlabeled node set $\mathcal{V}_{t}^{U}$ (blue). (b) The homophily (i.e., the fraction of intra-class pairs) of selected meta-path-based neighbor pairs $\mathcal{M}_{P}^{U}$ (orange) is much higher than that of the raw meta-path-based neighbor pair set $\mathcal{E}_{P}$ (blue).
@@
We also conduct an in-depth analysis to demonstrate the effectiveness of our reliable node selection strategy in RND and our intra-class meta-path-based neighbor estimation strategy in RMPD. Figure 4(a) presents the accuracy of raw unlabelled nodes $\mathcal{V}_{t}^{U}$, selected unlabelled reliable nodes $\mathcal{R}^{U}$, and all selected reliable nodes $\mathcal{R}$ with a selection proportion of $p=0.9$. Figure 4(b) displays the homophily of all meta-path-$P$-based neighbor pairs $\mathcal{E}_{P}$, selected unlabelled meta-path-$P$-based neighbor pairs $\mathcal{M}_{P}^{U}$, and all selected meta-path-$P$-based neighbor pairs $\mathcal{M}_{P}$. Here, homophily represents the fraction of intra-class neighbor pairs, with meta-paths MAM, QTQ, PAP, and PAP applied to TMDB, CroVal, ArXiv, and IGBs, respectively. We observed that the accuracy of selected unlabelled reliable nodes (orange) is significantly higher than that of raw unlabelled reliable nodes (blue), and the homophily of selected unlabelled meta-path-$P$-based neighbor pairs (orange) is also markedly higher than that of all meta-path-$P$-based neighbor pairs (blue). Furthermore, the accuracy of the final reliable nodes (green) used for distillation and the homophily of the final reliable intra-class meta-path-based neighbor pairs (green) used for distillation both exceed 0.9. These findings underscore the efficacy of our reliable node selection strategy in RND for identifying reliable nodes and our intra-class meta-path-based neighbor estimation strategy in RMPD for identifying intra-class neighbors.
TABLE IX: Ablation study on the effect of different meta-paths in RMPD.
Dataset w/ None w/ MP1 w/ MP2 w/ Both
TMDB 83.67±0.38 84.38±0.32 84.29±0.33 84.69±0.39
CroVal 88.05±0.12 88.37±0.06 88.76±0.04 89.05±0.10
IGB-549K-19 59.24±0.47 59.56±0.58 59.42±0.50 59.62±0.49
V-F How does each meta-path contribute to HG2M+? (RQ5)
In our main experiments, we use all meta-paths listed in Table I for RMPD. To better understand the contribution of each meta-path and examine possible redundancy, we conduct ablation studies on three representative datasets: TMDB, CroVal, and IGB-549K-19. Specifically, we compare four settings: (1) without using any meta-path (w/ None), (2) using only the first meta-path in Table I (w/ MP1, i.e., HG2M+ w/o RMPD), (3) using only the second meta-path (w/ MP2), and (4) using both (w/ Both, i.e., HG2M+). The results shown in Table IX demonstrate that each individual meta-path improves performance over the baseline without RMPD, and using both meta-paths yields the best results, confirming that each meta-path contributes unique, non-redundant information. Thus, meta-paths are complementary and beneficial to the overall model performance.
V-G How do HG2Ms perform with different teachers? (RQ6)
@@
V-H How do HG2Ms perform with noisy node features? (RQ7)
Considering that HG2Ms utilize only node features as input, they may be susceptible to feature noise and could underperform when the labels are unrelated to the node features. Therefore, we evaluate the robustness of HG2Ms with regards to different noise levels across TMDB, CroVal, and IGB-549K-19 datasets. Specifically, we introduce different levels of Gaussian noises to node features by replacing $\boldsymbol{X}_{t}$ with $(1-\alpha)\boldsymbol{X}_{t}+\alpha\epsilon$, where $\epsilon$ is independent isotropic Gaussian noise, and $\alpha\in[0,1]$ controls the noise level. As depicted in Figure 6, HG2Ms achieve comparable or improved performance compared to teacher HGNNs across different noise levels. Moreover, HG2Ms not only outperform MLPs but also exhibit slower performance degradation as $\alpha$ increases. When $\alpha$ approaches 1, the input features and node labels will become independent corresponding to the extreme case discussed in Section IV-C. In this scenario, HG2M+ continues to perform as well as HGNNs, while MLPs perform poorly.
V-I How do different hyperparameters affect HG2Ms? (RQ8)
V-I1 Inductive Split Rate